OpenStudy (vshiroky):

A sample was decomposed in the laboratory and found to have 38.67g C, 16.22g H, and 45.11g N. a)Find the molecular formula of this compound if the Formula mass is 62.12 g/mole. b)Determine how many H atoms would be in a 3.50g sample of this compound.

OpenStudy (anonymous):

To start you convert the masses of each element to moles. \[38.67 g /12.01 \frac{ g }{ mol C} = 3.22mol C\] \[16.22g / 1.01 \frac{ g }{ mol H} = 16.06mol H\] \[45.11g/14.01\frac{ g }{ mol N} = 3.22mol N\] You then can divide these by the lowest value, the moles of either Carbon or Nitrogen in this case. Round if needed. \[3.22mol C /3.22mol C = 1.00\] \[16.06mol H /3.22mol C = 4.99\approx 5.00\] \[3.22mol N /3.22mol C = 1.00\] This allows us to build the empirical formula. \[CNH_5\] We then compute the molar mass of the empirical formula. \[1mol C \times 12.01\frac{ g }{ mol C } + 1mol N \times \frac{ g }{ mol N} + 5mol H \times \frac{g}{mol H} = 31.07\] So we know we need to double the empirical formula to get the molecular weight specified. \[(CNH_5)_2\] or \[C_2N_2H_{10}\]

OpenStudy (anonymous):

To determine how many H atoms are in a sample, we need to find out how many moles that sample is, multiply that by the number of H atoms per molecule, and then convert that number from moles to atoms. \[3.50g / \frac{62.12g}{mol} = 0.0563mol\] \[0.0563mol \times \frac{10molH}{mol} = 0.563mol\] \[0.563molH \times 6.02 \times 10^{23} \frac{atoms}{mol} = 3.39 \times 10^{23} atomsH\]